# Maximum Likelihood Estimates of Parameters in Generalized Gamma Distribution with SeLF Algorithm

### v1

@@ 1.1 Introduction
The Gamma distribution is a commonly used probability distribution that has many applications in various fields, such as engineering, physics, and finance. It is a continuous probability distribution that is widely used to model waiting times, survival analysis, and lifetime data. However, there are situations where the Gamma distribution may not be an appropriate model due to its restrictive shape. In the paper [1], [1] have introduced a generalization of the Gamma distribution called the generalized Gamma distribution (GGD), which allows for more flexibility in the shape of the distribution. In this section, we discuss the properties and applications of this distribution and compare it to the (traditional) Gamma distribution.
@@ 1.2 Definition of Generalized Gamma Distribution
The generalized Gamma distribution is a distribution with the general formula of probability density function (PDF) given by:
@@ 1.4 Applications of Generalized Gamma Distribution
  • Lifetime Data: The generalized Gamma distribution can be used to model lifetime data, such as the time to failure of a component in a machine. For a more detailed discussion, please refer to [2]’s paper [2].
  • Wind Speed: The generalized Gamma distribution can be used to model wind speed data, which is important in the design of wind turbines. For a more detailed discussion, please refer to the paper [3] by [3].
  • Finance: The generalized Gamma distribution can be used to model the distribution of stock returns, which can help in the analysis of financial data. For a more detailed discussion, please refer to [4]’s paper [4].
@@ 1.5 Comparison with Gamma Distribution
The gamma distribution is a family of continuous probability distributions with two parameters including a shape parameter $\alpha$ and a scale parameter $\beta$. Its probability density function (PDF) is given by:
@@ 2.1 Introduction
The paper [5] by [5] has proposed the second-derivate lower-bound function (SeLF) algorithm, which is a novel maximization method based on the framework of minorization-maximization (MM) algorithms. The SeLF algorithm is a useful tool for calculating the maximum likelihood estimate (MLE) $\hat{\theta}$ of some parameter in a one-dimemsional log-likelihood function $l(\theta)$ (or maximize a more generalized target function). In order to introduce the SeLF algorithm more clearly, we have to spend some space introducing MM algorithms first, because the SeLF algorithm is based on the fundamental principles of MM algorithms.
@@ 2.2 Minorization-Maximization Algorithms
Minorization-Maximization (MM) algorithms are a family of iterative optimization algorithms that are commonly used in machine learning and statistical inference. These algorithms are particularly useful for solving optimization problems that involve log-concave likelihood functions or other functions that are difficult to optimize directly. In this subsection, we will provide a formal definition of MM algorithms and discuss some of their key properties and applications. [6] compiled the general principles of MM algorithms in their paper [6].
@@ 2.3.1 Definition
The paper [5] proposes a novel minorization-maximization (MM) method known as the second-derivative lower-bound function (SeLF) algorithm, which is a general approach for iteratively computing the maximum likelihood estimate (MLE) $\hat{\theta}$ of the parameter $\theta$ in a one-dimensional target function of interest, typically the marginal likelihood function $l(\theta)$.
@@
Since $h^{\prime\prime}(\theta|\theta^{(t)})\geq 0$ indicating that $h(\theta)$ is convex, by Taylor formula from analysis, we can see that for some specific $\theta^{*}$ between $\theta$ and $\theta^{(t)}$
@@ 2.3.2 Properties
Several important properties are described by the paper [5] as follows.
@@
Definition 3 (Weakly Stable Convergence[5]).
@@
Definition 4 (Strongly Stable Convergence[5]).
@@
The paper [5] established the following theorem:
Theorem 1 (Strongly Stable Convergence[5]).
@@
The proof of Theorem 1 in the paper [5] also guarantees the stability property.
@@ 3.1 Introduction
In [7]’s recent work [7], [7] introduce a new and versatile root-finding method, named the Upper-crossing/Solution (US) algorithm, which belongs to the class of non-bracketing (or open domain) methods. The US algorithm offers a general approach to iteratively determine the unique root $\theta^{*}$ of a non-linear equation $g(\theta)=0$. Each iteration of the US algorithm consists of two steps: an upper-crossing step (U-step) and a solution step (S-step). During the U-step, the algorithm seeks an upper-crossing function or a $U$-function, denoted by $U(\theta|\theta^{(t)})$, which takes a particular form depending on the $t$-th iteration $\theta^{(t)}$ of $\theta^{*}$, and is derived using a novel concept called the changing direction inequality. In the subsequent S-step, the US algorithm solves the simple $U$-equation, $U(\theta|\theta^{(t)})=0$, to obtain the explicit solution $\theta^{(t+1)}$.
@@ 3.2 Definition
Definition 5 (Changing Direction (CD) Inequalities[7]).
@@
Alternatively, we can simplify the assumption with CD inequalities as $g(\theta)\overset{{\rm sgn}(\theta-\theta^{*})}{\leq}0$ for all $\theta\in\Theta$. If $g(\theta)<0$ when $\theta<\theta^{*}$, then we can multiply both sides of the nonlinear equation $g(\theta)=0,\ \theta\in\Theta\subset\mathbb{R}$ by $-1$ to obtain a new equation $-g(\theta)=0$, which satisfies the above assumption.
@@
The conditions indicate that the $U(\theta|\theta^{(t)})$ function dominates the objective function $g(\theta)$ at $\theta=\theta^{(t)}$ for all $\theta\leq\theta^{(t)}$ and the $U(\theta|\theta^{(t)})$ function underestimates $g(\theta)$ at $\theta=\theta^{(t)}$ for all $\theta\geq\theta^{(t)}$. Using the definition of the $U$-function, we conclude that $g(\theta)$ is a $U$-function for 00 at $\theta=\theta^{*}$.
Some basic properties of $U$-functions are give by the paper [7] as follows.
Proposition 1 (Basic Properties of $U$-functions[7]).
@@
    If $U_{1}(\theta|\theta^{(t)})$ and $U_{2}(\theta|\theta^{(t)})$ are two $U$-functions for $g(\theta)$ at $\theta=\theta^{(t)}$, then $[U_{1}(\theta|\theta^{(t)})+U_{2}(\theta|\theta^{(t)})]/2$ is also a $U$-function for $g(\theta)$ at $\theta=\theta^{(t)}$;
@@
If a $U$-function can be found, one need only solve the simple $U$-equation: $U(\theta|\theta^{(t)})=0$ to obtain its solution $\theta^{(t+1)}$. This is denoted as
@@
Therefore, the US algorithm is a general principle for iteratively seeking the root $\theta^{*}$ with strongly stable converge. Each iteration consists of an upper-crossing step (U-step) and a solution step (S-step), where

### v2

@@ 1.1 Introduction
The Gamma distribution is a commonly used probability distribution that has many applications in various fields, such as engineering, physics, and finance. It is a continuous probability distribution that is widely used to model waiting times, survival analysis, and lifetime data. However, there are situations where the Gamma distribution may not be an appropriate model due to its restrictive shape. In the paper [stacy1962generalization], [stacy1962generalization] has introduced a generalization of the Gamma distribution called the generalized Gamma distribution (GGD), which allows for more flexibility in the shape of the distribution. In this section, we discuss the properties and applications of this distribution and compare it to the (traditional) Gamma distribution.
@@ 1.2 Definition of Generalized Gamma Distribution
The generalized Gamma distribution is a distribution with the general formula of the probability density function (PDF) given by:
@@ 1.4 Applications of Generalized Gamma Distribution
  • Lifetime Data: The generalized Gamma distribution can be used to model lifetime data, such as the time to failure of a component in a machine. For a more detailed discussion, please refer to [shanker2016modeling]’s paper [shanker2016modeling].
  • Wind Speed: The generalized Gamma distribution can be used to model wind speed data, which is important in the design of wind turbines. For a more detailed discussion, please refer to the paper [mert2015statistical] by [mert2015statistical].
  • Finance: The generalized Gamma distribution can be used to model the distribution of stock returns, which can help in the analysis of financial data. For a more detailed discussion, please refer to [gomes2006four]’s paper [gomes2006four].
@@ 1.5 Comparison with Gamma Distribution
The gamma distribution is a family of continuous probability distributions with two parameters, including a shape parameter $\alpha$ and a scale parameter $\beta$. Its probability density function (PDF) is given by:
@@ 2.1 Introduction
The paper [tian2025second] by [tian2025second] has proposed the second-derivative lower-bound function (SeLF) algorithm, which is a novel maximization method based on the framework of minorization-maximization (MM) algorithms. The SeLF algorithm is a useful tool for calculating the maximum likelihood estimate (MLE) $\hat{\theta}$ of some parameter in a one-dimensional log-likelihood function $l(\theta)$ (or maximize a more generalized target function). In order to introduce the SeLF algorithm more clearly, we have to spend some space introducing MM algorithms first, because the SeLF algorithm is based on the fundamental principles of MM algorithms.
@@ 2.2 Minorization-Maximization Algorithms
Minorization-Maximization (MM) algorithms are a family of iterative optimization algorithms that are commonly used in machine learning and statistical inference. These algorithms are particularly useful for solving optimization problems that involve log-concave likelihood functions or other functions that are difficult to optimize directly. In this subsection, we will provide a formal definition of MM algorithms and discuss some of their key properties and applications. [hunter2004tutorial] compiled the general principles of MM algorithms in their paper [hunter2004tutorial].
@@ 2.3.1 Definition
The paper [tian2025second] proposes a novel minorization-maximization (MM) method known as the second-derivative lower-bound function (SeLF) algorithm, which is a general approach for iteratively computing the maximum likelihood estimate (MLE) $\hat{\theta}$ of the parameter $\theta$ in a one-dimensional target function of interest, typically the marginal likelihood function $l(\theta)$.
@@
Since $h^{\prime\prime}(\theta|\theta^{(t)})\geq 0$, indicating that $h(\theta)$ is convex, by the Taylor formula from analysis, we can see that for some specific $\theta^{*}$ between $\theta$ and $\theta^{(t)}$
@@ 2.3.2 Properties
Several important properties are described by the paper [tian2025second] as follows.
@@
Definition 3 (Weakly Stable Convergence[tian2025second]).
@@
Definition 4 (Strongly Stable Convergence[tian2025second]).
@@
The paper [tian2025second] established the following theorem:
Theorem 1 (Strongly Stable Convergence[tian2025second]).
@@
The proof of Theorem 1 in the paper [tian2025second] also guarantees the stability property.
@@ 3.1 Introduction
In [li2022upper]’s recent work [li2022upper], [li2022upper] introduces a new and versatile root-finding method, named the Upper-crossing/Solution (US) algorithm, which belongs to the class of non-bracketing (or open domain) methods. The US algorithm offers a general approach to iteratively determine the unique root $\theta^{*}$ of a non-linear equation $g(\theta)=0$. Each iteration of the US algorithm consists of two steps: an upper-crossing step (U-step) and a solution step (S-step). During the U-step, the algorithm seeks an upper-crossing function or a $U$-function, denoted by $U(\theta|\theta^{(t)})$, which takes a particular form depending on the $t$-th iteration $\theta^{(t)}$ of $\theta^{*}$, and is derived using a novel concept called the changing direction inequality. In the subsequent S-step, the US algorithm solves the simple $U$-equation, $U(\theta|\theta^{(t)})=0$, to obtain the explicit solution $\theta^{(t+1)}$.
@@ 3.2 Definition
Definition 5 (Changing Direction (CD) Inequalities[li2022upper]).
@@
Alternatively, we can simplify the assumption with CD inequalities as $g(\theta)\overset{{\rm sgn}(\theta-\theta^{*})}{\leq}0$ for all $\theta\in\Theta$. If $g(\theta)<0$ when $\theta<\theta^{*}$, then we can multiply both sides of the nonlinear equation $g(\theta)=0,\ \theta\in\Theta\subset\mathbb{R}$ by 1-1 to obtain a new equation $g(\theta)=0$, which satisfies the above assumption.
@@
The conditions indicate that the $U(\theta|\theta^{(t)})$ function dominates the objective function $g(\theta)$ at $\theta=\theta^{(t)}$ for all $\theta\leq\theta^{(t)}$ and the $U(\theta|\theta^{(t)})$ function underestimates $g(\theta)$ at $\theta=\theta^{(t)}$ for all $\theta\geq\theta^{(t)}$. Using the definition of the $U$-function, we conclude that $g(\theta)$ is a $U$-function for 0 at $\theta=\theta^{*}$.
Some basic properties of $U$-functions are given by the paper [li2022upper] as follows.
Proposition 1 (Basic Properties of $U$-functions[li2022upper]).
@@
    If $U_{1}(\theta|\theta^{(t)})$ and $U_{2}(\theta|\theta^{(t)})$ are two $U$-functions for $g(\theta)$ at $\theta=\theta^{(t)}$, then $U_{1}(\theta|\theta^{(t)})+U_{2}(\theta|\theta^{(t)})]/2$ is also a $U$-function for $g(\theta)$ at $\theta=\theta^{(t)}$;
@@
If a $U$-function can be found, one needs only solve the simple $U$-equation: $U(\theta|\theta^{(t)})=0$ to obtain its solution $\theta^{(t+1)}$. This is denoted as
@@
Therefore, the US algorithm is a general principle for iteratively seeking the root $\theta^{*}$ with strongly stable convergence. Each iteration consists of an upper-crossing step (U-step) and a solution step (S-step), where